A NEW POTENTIAL ENERGY SURFACE FOR N_{2}O–He, AND PIMC SIMULATIONS PROBING IN-FRARED SPECTRA AND SUPERFLUIDITY

\underline{LECHENG WANG}, DAIQIAN XIE, School of Chemistry and Chemical Engineering, Nanjing University, No.22 Hankou Road, Nanjing, Jiangsu, China, 210093 ; HUI LI, ROBERT J. LE ROY, and PIERRE-NICHOLAS ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

High resolution infrared spectra of an N_{2}O molecule seeded in ultra-cold helium custers have been reported by two groups.^{a}^{,}^{b} Early simulations of N_{2}O–(He)_{n} clusters^{c}^{,}^{d} were performed using potential energy surfaces (PESs) for which the dopant N_{2}O molecule was frozen at its equilibrium geometry. Since the evolution of the shift of the ν_{3} band-origin of N_{2}O provides a key link to bridge the gap between micro and macro world, a new 3D PES was generated which incorporated the asymmetric-stretch Q_{3} vibrational motion of the N_{2}O.^{e} Bosonic PIMC simulations^{f} based on this surface were then used to study rotational dynamics, and energetic and superfluid properties of N_{2}O–(He)_{n} clusters. The evolution of the calculated shifts agree reasonably with the experimental results, but some quantitative discrepancies remain. To address this problem, a new four-dimensional N_{2}O–He PES has now been obtained which also takes account of the change in the average value of the Q_{1} symmetric-stretch coordinate on excitation of ν_{3}. It has been fitted to a generalized MLR functional form^{g} which imposes better long-range behaviour. Results obtained using this new surface will be presented.

a Y. Xu, W. Jager, J. Tang and A.R.W. McKellar, Phys. Rev. Lett. 91, 163401 (2003).

b A.R.W. McKellar, J. Chem. Phys. 127, 044315 (2007).

c S. Moroni, N. Blinov and P-N. Roy, J. Chem. Phys. 121, 3577 (2004).

d F. Paesani and K.B. Whaley, J. Chem. Phys. 121, 5293 (2004).

e Y.Z. Zhou, D. Xie and D.H. Zhang, J. Chem. Phys. 124, 144317(2006).

f H. Li, N. Blinov, P.-N. Roy and R.J. Le Roy, J. Chem. Phys. 130, 144305(2009).

gH. Li and R.J. Le Roy, Phys. Chem. Chem. Phys. 10, 04128 (2008).